### v1

@@ 4.1 Unconstrained convex quadratic minimization
4.1.1 Extension: $\ell_{1}$-regularized convex quadratic minimization
[…]
4.2 Minimization of a polynomial over a rectangle
In this subsection, we consider the minimization of a polynomial function of degree $N\in\mathbb{N}$ over a hyper-rectangle in $\mathbb{R}^{p}$:
$\begin{array}[]{|ll}\underset{\bm{\lambda}\in\mathbb{R}^{p}}{\mbox{minimize}}&F(\bm{\lambda})=\sum\limits_{i\in[I]}a_{i}\prod\limits_{j\in[p]}\lambda_{j}^{n_{i,j}}\\ \mbox{subject to}&\bm{l}\leq\bm{\lambda}\leq\bm{u},\end{array}$ (29)
where $a_{i}\in\mathbb{R}$ for $i\in[I]$, $n_{i,j}\in\{0\}\cup[N]$ for $i\in[I],j\in[p]$, and $\bm{l},\bm{u}\in\mathbb{R}^{p}$. For example, the dual form of Support Vector Classification for a bias-free decision function can be cast in this form with a convex quadratic objective. We can see that for a constant $K$ such that $K>\max\limits_{\bm{\lambda}\in[\bm{l},\bm{u}]}F(\bm{\lambda})$, (29) can be equivalently rewritten as
$\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}&f(\bm{\theta})=-\ln\Big{(}K-F\big{(}\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{l}\big{)}\Big{)}\\ \mbox{subject to}&\bm{0}\leq\bm{\theta}\leq\bm{1},\end{array}$ (32)
where $\bm{1}:=(1,...,1)^{\top}$. Note that there exists an upper bound $K$ because $F$ is continuous and the feasible region is compact. See Appendix A.6 for an example of how to obtain such $K$.
To apply the EM scheme to (32), assume that $\bm{X}=(X_{1},...,X_{p})$ is a vector of $p$ independent binomial random variables, $X_{j}\sim\mathrm{Bin}(m_{j},\theta_{j})$, with $m_{j}:=\max\{n_{i,j}\,|\,i\in[I]\},~{}j\in[p]$.$\bm{X}$ is then of the exponential family (11) since its probability mass on $\bm{x}\in\Xi=\prod\limits_{j\in[p]}(\{0\}\cup[m_{j}])$ is given by
$p_{\bm{x}}(\bm{\theta})$ $=\prod_{j\in[p]}\frac{m_{j}!}{x_{j}!(m_{j}-x_{j})!}\theta_{j}^{x_{j}}(1-\theta_{j})^{m_{j}-x_{j}}$ (33)
$=\underbrace{\prod_{j\in[p]}\frac{m_{j}!}{x_{j}!(m_{j}-x_{j})!}}_{=\beta(\bm{x})}\exp\Big{(}\underbrace{\sum_{j\in[p]}x_{j}\ln\frac{\theta_{j}}{1-\theta_{j}}}_{=\bm{\eta}(\bm{\theta})^{\top}\bm{\gamma}(\bm{x})}+\underbrace{\sum_{j\in[p]}m_{j}(1-\theta_{j})}_{-A(\bm{\theta})}\Big{)}.$
To derive a function $G(\cdot)$ that satisfies (4) for (32) under (33), note first that there exist constants $\tilde{a}_{i},\,i\in[I]$, and $\tilde{b}$ such that
$F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{l})$ $\equiv\sum_{i\in[I]}a_{i}\prod_{j\in[p]}\big{(}(u_{j}-l_{j})\theta_{j}+l_{j}\big{)}^{n_{i,j}}=\sum_{i\in[I]}\tilde{a}_{i}\prod_{j\in[p]}\theta_{j}^{n_{i,j}}+\tilde{b}.$
Proposition 4.4.
If $G$ is defined by
$G(\bm{x})$ $=K-\Big{(}\sum_{i\in[I]}\tilde{a}_{i}\prod_{j\in[p]}g_{i,j}(x_{j})+\tilde{b}\Big{)},$ (34)
where $g_{i,j}(\cdot)$ is defined with $\overline{m}_{j}\in\mathbb{N}$ such that $\overline{m}_{j}\geq m_{j}\equiv\max\{n_{i,j}\,|\,i\in[I]\},$ as
$g_{i,j}(x_{j}):=$$\begin{cases}\displaystyle\frac{x_{j}(x_{j}-1)\cdots(x_{j}-n_{i,j}+1)}{\overline{m}_{j}(\overline{m}_{j}-1)\cdots(\overline{m}_{j}-n_{i,j}+1)},&(n_{i,j}\in[N]),\\ 1,&(n_{i,j}=0),\end{cases}$
the condition (4) for (32) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (33) and $\Theta=[\bm{0},\bm{1}]$.
See Appendix A.5 for the proof. While the above statement holds for any $\overline{m}_{j}\geq m_{j}$, we will only use $m_{j}$ for $\overline{m}_{j}$, so that the longest step size will be taken.
The EM algorithm iteratively minimizes the surrogate function
$u(\bm{\theta}|\bar{\bm{\theta}})$ $=-\sum\limits_{j\in[p]}\left\{\mathbb{E}_{q_{\bm{x}}(\bar{\theta}_{j})}[X_{j}]\ln\frac{\theta_{j}}{\bar{\theta}_{j}}+(m_{j}-\mathbb{E}_{q_{\bm{x}}(\bar{\theta}_{j})}[X_{j}])\ln\frac{1-\theta_{j}}{1-\bar{\theta}_{j}}\right\}+f(\bar{\bm{\theta}}),$
@@ 4.2 Minimization of a polynomial over a rectangle
If the binomial distribution (33) is employed for Algorithm 1 to solve (32), the update formula (19) results in
$\theta^{(t)}_{j}\leftarrow\theta^{(t-1)}_{j}-\frac{\theta^{(t-1)}_{j}(1-\theta^{(t-1)}_{j})}{m_{j}}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)}),\quad j\in[p].$ (35)
@@
Corollary 4.5 reveals that minimizing the surrogate (or (16)) results in a component-wise gradient descent (35) whose step size is endogenously determined at each iteration as $\theta_{j}^{(t-1)}\big{(}1-\theta_{j}^{(t-1)}\big{)}/m_{j},~{}j\in[p]$. Note also that if we start the algorithm from an interior point of the rectangle, the generated sequence $\{\bm{\theta}^{(t)}\}_{t\geq 0}$ will remain in the interior, $(0,1)^{p}$, which implies that the above algorithm is an interior point method. Algorithm 3 shows the full description of the procedure.
Algorithm 3 EM algorithm for minimizing the polynomial objective over a rectangle (32)
1:Let $K>\max_{\bm{\lambda}\in[\bm{l},\bm{u}]}F(\bm{\lambda})$ and $\bm{\theta}^{0}\in(0,1)^{p}$ , and set $t\leftarrow 1$.
@@
3:     Update the solution by (35) and $t\leftarrow t+1$.
@@
4.3 Minimization of a polynomial over a unit simplex
As the third example, this section considers the minimization of the same objective function as (29) over a unit simplex.
$({\rm Poly.Simplex}_{0})\quad\begin{array}[]{|ll}\underset{\bm{\lambda}\in\mathbb{R}^{p}}{\mbox{minimize}}&F(\bm{\lambda}):=\displaystyle\sum\limits_{i\in[I]}a_{i}\prod\limits_{j\in[p]}\lambda_{j}^{n_{i,j}}\\ \mbox{subject to}&\bm{1}^{\top}\bm{\lambda}=1,~{}\bm{\lambda}\geq\bm{0}.\\ \end{array}$ (38)
(38) encompasses, for example, a portfolio selection problem in which the objective function consisting of the weighted sum of higher order moments of asset returns is optimized without a short sale. Similar to (32), for a sufficiently large $K$, rewrite (38) as
$({\rm Poly.Simplex^{\prime}})\quad\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{p-1}}{\mbox{minimize}}&f(\bm{\theta}):=-\ln\big{(}K-F((\bm{I},-\bm{1})^{\top}\bm{\theta}+\bm{e}_{p})\big{)}\\ \mbox{subject to}&\bm{1}^{\top}\bm{\theta}\leq 1,~{}\bm{\theta}\geq\bm{0},\end{array}$ (41)
where $\bm{I}\in\mathbb{R}^{p-1\times p-1},\bm{1}\in\mathbb{R}^{p-1},\bm{e}_{p}=(0,\ldots,0,1)^{\top}\in\mathbb{R}^{p}$. Note that $\lambda_{k}=\theta_{k}$ for $k\in[p-1]$; $\lambda_{p}=1-\bm{1}^{\top}\bm{\theta}$. See Appendix A.9 for an example how to get $K$.
@@ 4.3 Minimization of a polynomial over a unit simplex
$p_{\bm{x}}(\bm{\theta})$ $=\frac{m!}{x_{1}!\cdots x_{p-1}!\big{(}m-\sum_{j\in[p-1]}x_{j}\big{)}!}\theta_{1}^{x_{1}}\cdots\theta_{p-1}^{x_{p-1}}\big{(}1-\sum\limits_{j\in[p-1]}\theta_{j}\big{)}^{m-\sum_{j\in[p-1]}x_{j}}$ (42)
$=\underbrace{\frac{m!}{x_{1}!\cdots x_{p-1}!\big{(}m-\sum_{j\in[p-1]}x_{j}\big{)}!}}_{=\beta(\bm{x})}\exp\Big{(}\underbrace{\sum_{j\in[p-1]}x_{j}\ln\frac{\theta_{j}}{1-\sum_{j\in[p-1]}\theta_{j}}}_{=\bm{\gamma}(\bm{x})^{\top}\bm{\eta}(\bm{\theta})}\underbrace{+m\ln\big{(}1-\sum\limits_{j\in[p-1]}\theta_{j}\big{)}}_{=-A(\bm{\theta})}\Big{)}.$
@@
If $G$ is defined by
$G(\bm{x})$ $:=K-\sum_{i\in[I]}a_{i}g_{i}(\bm{x}),$
where $g_{i}(\cdot)$ is defined with $\overline{m}\in\mathbb{N}$ such that $\overline{m}\geq m\equiv\max\big{\{}\sum\limits_{j\in[p]}n_{i,j}\big{|}\,i\in[I]\big{\}}$, as
$g_{i}(\bm{x}):=$$\left\{\begin{array}[]{ll}\displaystyle\frac{\prod\limits_{j\in[p]}x_{j}(x_{j}-1)\cdots(x_{j}-n_{i,j}+1)}{\overline{m}(\overline{m}-1)\cdots\Big{(}\overline{m}-\sum\limits_{j\in[p]}n_{i,j}+1\Big{)}},&(n_{i,j}\in[N]),\\ 1,&(n_{i,j}=0),\end{array}\right.$
where $\bm{x}=(x_{1},\ldots,x_{p-1})^{\top}$ and $x_{p}=m-\bm{1}^{\top}\bm{x}$, the condition (4) for (41) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (42) and $\Theta=\{\bm{\theta}\in\mathbb{R}_{+}^{p-1}|\bm{1}^{\top}\bm{\theta}\leq 1\}$.
Similar to the optimization over a rectangle, the surrogate function is given as
$u(\bm{\theta}|\bar{\bm{\theta}})=-\sum\limits_{j\in[p-1]}\left\{\mathbb{E}_{q_{\bm{x}}(\bar{\theta}_{j})}[X_{j}]\ln\frac{\theta_{j}}{\bar{\theta}_{j}}+(m-\mathbb{E}_{q_{\bm{x}}(\bar{\theta}_{j})}[X_{j}])\ln\frac{1-\sum_{h\in[p-1]}\theta_{h}}{1-\sum_{h\in[p-1]}\bar{\theta}_{h}}\right\}+f(\bar{\bm{\theta}}),$
and its minimization can be reduced to a natural gradient descent as follows.
@@
If the multinomial distribution (42) is employed for Algorithm 1 to solve (41), the update formula (19) results in
$\begin{split}V^{(t-1)}&\leftarrow\sum_{j\in[p-1]}\theta_{j}^{(t-1)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)}),\\ \theta_{j}^{(t)}&\leftarrow\theta_{j}^{(t-1)}-\frac{\theta^{(t-1)}_{j}}{m}\Big{(}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)})-V^{(t-1)}\Big{)},\quad j\in[p-1].\end{split}$ (43)
Similar to (35), the update formula (43) is component-wise and no line search is involved. See Appendix A.10 for the proof. Algorithm 4 shows the full description of the procedure.
Algorithm 4 EM algorithm for minimizing the polynomial objective over a simplex (41)
1:Let $K>\max\{F(\bm{\theta})|\bm{1}^{\top}\bm{\theta}\leq 1,\bm{\theta}\geq\bm{0}\}$ and $\bm{\theta}^{0}\in\{\bm{\theta}\in\mathbb{R}_{++}^{p-1}|\bm{1}^{\top}\bm{\theta}\leq 1\}$, and set $t\leftarrow 1$.
@@
3:     Update the solution by (35) and $t\leftarrow t+1$.
@@
4.4 Global convergence
Global convergence of Algorithms 3 and 4 can be shown in a unified manner. Assume that the feasible set of (1) is represented by a system of inequalities as $\Theta=\{\bm{\theta}\in\mathbb{R}^{p}|g_{i}(\bm{\theta})\leq 0,i\in[m]\}$ and has an interior point, i.e., $\exists\bm{\theta}$ such that $g_{i}(\bm{\theta})<0,i\in[m]$:
$\begin{array}[]{|lll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}&f(\bm{\theta})\\ \mbox{subject to}{}{}&g_{i}(\bm{\theta})\leq 0,\quad i\in[m],\end{array}$ (46)

### v2

@@ 4.1. Unconstrained convex quadratic minimization
Let $\overline{\lambda}_{\bm{S}},\underline{\lambda}_{\bm{S}}$ denote the largest and smallest eigenvalues of a real symmetric matrix $\bm{S}$, respectively. Suppose the same assumptions as in Corollary 4.2. (i) We have $f(\bm{\theta}^{(t-1)})-f(\bm{\theta}^{(t)})\geq\|\bm{\theta}^{(t)}-\bm{\theta}^{(t-1)}\|^{2}/(2\overline{\lambda}_{\bm{\Sigma}}),$ which implies $\min_{t\in[T]}\|\bm{\theta}^{(t)}-\bm{\theta}^{(t-1)}\|^{2}\leq 2\overline{\lambda}_{\bm{\Sigma}}\big(f(\bm{\theta}^{(0)})-f(\bm{\theta}^{*})\big)/T.$ (ii) We have $f(\bm{\theta}^{(t-1)})-f(\bm{\theta}^{(t)})\geq(\underline{\lambda}_{\bm{\Sigma}}/2)\|\nabla f(\bm{\theta}^{(t-1)})\|^{2},$ which implies $\min_{t\in[T]}\|\nabla f(\bm{\theta}^{(t-1)})\|^{2}\leq(2/\underline{\lambda}_{\bm{\Sigma}})\big(f(\bm{\theta}^{(0)})-f(\bm{\theta}^{*})\big)/T.$ (iii) In addition, if $\underline{\lambda}_{\bm{Q}}>0$, we have $f(\bm{\theta}^{(T)})-f(\bm{\theta}^{*})\leq\left(1-\underline{\lambda}_{\bm{Q}}\underline{\lambda}_{\bm{\Sigma}}\right)^{T}\big(f(\bm{\theta}^{(0)})-f(\bm{\theta}^{*})\big).$
4.1.1. Extension: $\ell_{1}$-regularized convex quadratic minimization
With $f(\bm{\theta})=\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta}$ and $\rho(\bm{\theta})=\|\bm{\theta}\|_{1}:=\sum_{j\in[p]}|\theta_{j}|$, the composite optimization (2.10) becomes the $\ell_{1}$-regularized convex quadratic minimization:
$\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}~\Phi(\bm{\theta})=\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta}+\|\bm{\theta}\|_{1}.$
With the diagonal matrix $\bm{\Sigma}=\mathrm{diag}(\sigma_{1},...,\sigma_{p})$, the surrogate function becomes separable as $U(\bm{\theta}|\bar{\bm{\theta}})=u(\bm{\theta}|\bar{\bm{\theta}})+\|\bm{\theta}\|_{1}=\sum_{j\in[p]}\big((\theta_{j}-\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[X_{j}])^{2}/(2\sigma_{j})+|\theta_{j}|\big)+\mbox{``constant,''}$ and it can be minimized via the so-called soft-thresholding operation, i.e., for $\mathbb{E}_{q_{\bm{x}}(\bar{\bm{\theta}})}[X_{j}]=\bar{\theta}_{j}-\sigma_{j}\sum_{h\in[p]}(q_{jh}\bar{\theta}_{h}+b_{j}),~j\in[p],$ the coordinate-wise update formula is derived as $\theta_{j}^{(t)}\leftarrow{\rm soft}_{\sigma_{j}}(\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[X_{j}])$ where ${\rm soft}_{a}(z):=\arg\min_{x}\left\{a|x|+\frac{1}{2}(x-z)^{2}\right\}=z+a$ if $z\leq-a$0 if $a\leq z\leq a$zaz-a if $a\leq z$, defines the soft-thresholding with parameter $a\in\mathbb{R}$. This procedure is analogous to ISTA (Iterative Soft-Thresholding Algorithm) for the $\ell_{1}$-regularized convex quadratic minimization, yet it differs from the vanilla ISTA in that it permits the step size to vary for each coordinate. Also, while ISTA has famous applications in image restoration and reconstruction, and EM algorithms associated with normal distribution exist (e.g., [15]), the role of normal distribution in the above treatment has little to do with that in such statistical contexts.
4.2. Minimization of a polynomial over a rectangle
In this subsection, we consider the minimization of a polynomial function over a hyper-rectangle in $\mathbb{R}^{p}$:
$\begin{array}[]{|ll}\underset{\bm{\lambda}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&F(\bm{\lambda}):=\sum\limits_{i\in[I]}a_{i}\prod\limits_{j\in[p]}\lambda_{j}^{n_{i,j}}\\ \mbox{subject to}&\bm{l}\leq\bm{\lambda}\leq\bm{u},\end{array}$ (4.10)
where $a_{i}\in\mathbb{R}$ for $i\in[I]$, $n_{i,j}\in\{0\}\cup\mathbb{N}$ for $i\in[I],j\in[p]$, and $\bm{l},\bm{u}\in\mathbb{R}^{p}$ satisfy $l_{j}<u_{j}$ for each $j$. For example, the dual form of Support Vector Classification for a bias-free decision function can be cast in this form with a convex quadratic objective. To establish a binomial distribution-based EM algorithm for this optimization problem, we first note that by the change of variables $\theta_{j}:=(\lambda_{j}-l_{j})/(u_{j}-l_{j})$, the problem (4.10) can be rewritten as
$\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&\sum\limits_{i\in[I]}\tilde{a}_{i}\prod\limits_{j\in[p]}\theta_{j}^{n_{i,j}}+\tilde{b}\quad(=F\big(\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{l}\big))\\ \mbox{subject to}&\bm{0}\leq\bm{\theta}\leq\bm{1},\end{array}$ (4.13)
for some $\tilde{a}_{i}\in\mathbb{R},i\in[I]$, and $\tilde{b}\in\mathbb{R}$ such that
$F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{l})$ $=\sum_{i\in[I]}a_{i}\prod_{j\in[p]}\big((u_{j}-l_{j})\theta_{j}+l_{j}\big)^{n_{i,j}}=\sum_{i\in[I]}\tilde{a}_{i}\prod_{j\in[p]}\theta_{j}^{n_{i,j}}+\tilde{b}$.
Note also that there exists an upper bound $K$ of the objective value of (4.10) and (4.13) because the objective function is continuous and the feasible region is compact. With such an upper bound $K$, i.e., $K>\max\{F(\bm{\lambda})|\bm{\lambda}\in[\bm{l},\bm{u}]\}(=\max\{F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{l})|\bm{\theta}\in[\bm{0},\bm{1}]\})$, (4.13) can be equivalently rewritten as
$\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&f(\bm{\theta})=-\ln\Big(K-F\big(\mathrm{diag}(\bm{u}-\bm{l})\bm{\theta}+\bm{l}\big)\Big)\\ \mbox{subject to}&\bm{0}\leq\bm{\theta}\leq\bm{1}.\end{array}$ (4.16)
See Appendix A.6 for an example of how to obtain such $K$.
To apply the EM scheme to (4.16), we consider a vector $\bm{X}=(X_{1},...,X_{p})$ of $p$ independent binomial random variables, $X_{j}\sim\mathrm{Bin}(m_{j},\theta_{j})$, with $m_{j}\geq\max\{n_{i,j}\,|\,i\in[I]\},~j\in[p]$, i.e., $m_{j}$ must be no less than the highest degree of variable $\theta_{j}$. $\bm{X}$ is then of the exponential family (3.1) since its probability mass on $\bm{x}\in\Xi=\prod_{j\in[p]}(\{0\}\cup[m_{j}])$ is given by
$p_{\bm{x}}(\bm{\theta})$ $=\prod_{j\in[p]}\frac{m_{j}!}{x_{j}!(m_{j}-x_{j})!}\theta_{j}^{x_{j}}(1-\theta_{j})^{m_{j}-x_{j}}$ (4.17)
$=\underbrace{\prod_{j\in[p]}\frac{m_{j}!}{x_{j}!(m_{j}-x_{j})!}}_{=\beta(\bm{x})}\exp\Big(\underbrace{\sum_{j\in[p]}x_{j}\ln\frac{\theta_{j}}{1-\theta_{j}}}_{=\bm{\gamma}(\bm{x})^{\top}\bm{\eta}(\bm{\theta})}+\underbrace{\sum_{j\in[p]}m_{j}\ln(1-\theta_{j})}_{=-A(\bm{\theta})}\Big).$
The next proposition shows a function $G(\cdot)$ that satisfies (2.4) for (4.16) under (4.17).
Proposition 4.4.
Let $m_{j}\geq\max\{n_{i,j}\,|\,i\in[I]\}$ for $j\in[p]$. If $G$ is defined by
@@ 4.2. Minimization of a polynomial over a rectangle
where, corresponding to $n_{i,j}$, $g_{i,j}(\cdot)$ is defined with $m_{j}$ as
$g_{i,j}(x_{j}):=$ $\begin{cases}\displaystyle\frac{x_{j}(x_{j}-1)\cdots(x_{j}-n_{i,j}+1)}{m_{j}(m_{j}-1)\cdots(m_{j}-n_{i,j}+1)},&\mbox{if }n_{i,j}>0,\\ 1,&\mbox{if }n_{i,j}=0,\end{cases}$
the condition (2.4) for (4.16) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (4.17) and $\Theta=[\bm{0},\bm{1}]$.
See Appendix A.5 for the proof.
Applying Proposition 3.1 (b), we see that the EM algorithm iteratively minimizes the surrogate function
$u(\bm{\theta}|\bm{\theta}^{(t-1)})$
$=-\sum\limits_{j\in[p]}\Big\{\mathbb{E}_{q_{\bm{x}}(\theta^{(t-1)}_{j})}[X_{j}]\ln\frac{\theta_{j}}{\theta^{(t-1)}_{j}}+(m_{j}-\mathbb{E}_{q_{\bm{x}}(\theta^{(t-1)}_{j})}[X_{j}])\ln\frac{1-\theta_{j}}{1-\theta^{(t-1)}_{j}}\Big\}+f(\bm{\theta}^{(t-1)}),$
@@
If the binomial distribution (4.17) is employed for Algorithm 1 to solve (4.16), the update formula (3.9) results in
$\theta^{(t)}_{j}\leftarrow\theta^{(t-1)}_{j}-\frac{\theta^{(t-1)}_{j}(1-\theta^{(t-1)}_{j})}{m_{j}}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)}),\quad j\in[p].$ (4.19)
@@
Corollary 4.5 reveals that minimizing the surrogate function (or (3.6)) for (4.16) results in a component-wise gradient descent (4.19) whose step size is endogenously determined at each iteration as $\theta_{j}^{(t-1)}\big(1-\theta_{j}^{(t-1)}\big)/m_{j},~j\in[p]$. Given that $m_{j}$ may take any number satisfying $m_{j}\geq\max\{n_{i,j}|i\in[I]\}$, the use of $m_{j}=\max\{n_{i,j}|i\in[I]\}$ yields the largest step size. Note also that if we start the algorithm from an interior point of the rectangle, the generated sequence $\{\bm{\theta}^{(t)}\}_{t\geq 0}$ will remain in the interior, $(0,1)^{p}$, which implies that the above algorithm is an interior point method. Algorithm 3 shows the full description of the procedure.
Algorithm 3 EM algorithm for minimizing a polynomial objective over a rectangle (4.16)
1:Let $K>\max_{\bm{\lambda}\in[\bm{l},\bm{u}]}F(\bm{\lambda})$ and $\bm{\theta}^{(0)}\in(0,1)^{p}$ , and set $t\leftarrow 1$.
@@
3:  Update the solution by (4.19) and $t\leftarrow t+1$.
@@
4.3. Minimization of a polynomial over a unit simplex
As the third example, this section considers the minimization of the same objective function as (4.10) over a unit simplex.
@@ 4.3. Minimization of a polynomial over a unit simplex
(4.22) encompasses, for example, a portfolio selection problem in which the objective function consisting of the weighted sum of higher order moments of asset returns is optimized without a short sale. To establish a multinomial distribution-based EM algorithm for (4.22), deleting the $p$-th variable $\lambda_{p}$ by using the relation $\lambda_{p}=1-\sum_{j\in[p-1]}\lambda_{j}$ and replacing $\lambda_{j}$ with $\theta_{j}$, $j\in[p-1]$, we get the equivalent formulation as
[…]
Similarly to (4.16), with an upper bound $K$ of the objective value, the problem (4.25) can be equivalently rewritten as
$\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{p-1}}{\mbox{minimize}}\quad&f(\bm{\theta})=-\ln\big(K-F((\bm{I},-\bm{1})^{\top}\bm{\theta}+\bm{e}_{p})\big)\\ \mbox{subject to}&\bm{1}^{\top}\bm{\theta}\leq 1,~\bm{\theta}\geq\bm{0},\end{array}$ (4.28)
where the size of $\bm{I}$ and $\bm{1}$ is p1p-1, i.e., $\bm{I}\in\mathbb{R}^{(p-1)\times(p-1)},\bm{1}\in\mathbb{R}^{p-1}$, and $\bm{e}_{p}$ denotes the unit vector whose $p$-th component is 1, i.e., $\bm{e}_{p}=(0,\ldots,0,1)^{\top}\in\mathbb{R}^{p}$. See Appendix A.9 for an example how to get $K$.
To apply the EM scheme to (4.28), consider the (p1)(p-1)-dimensional random variables $\bm{X}=(X_{1},\ldots,X_{p-1})$ that follows a multinomial distribution, $\bm{X}\sim\mathrm{Multi}(m,\bm{\theta})$, with $m\geq\max\{\sum_{j\in[p]}n_{i,j}|i\in[I]\}$. It is of the exponential family since its mass on $\bm{x}\in\Xi=\{\bm{x}\in(\{0\}\cup\mathbb{N})^{p-1}\,|\,\bm{1}^{\top}\bm{x}\leq m\}$ is given by
$p_{\bm{x}}(\bm{\theta})$ $=\frac{m!}{x_{1}!\cdots x_{p-1}!\big(m-\bm{1}^{\top}\bm{x}\big)!}\prod_{j\in[p-1]}\theta_{j}^{x_{j}}{\big(1-\bm{1}^{\top}\bm{\theta}\big)^{m-\bm{1}^{\top}\bm{x}}}$ (4.29)
$=\underbrace{\frac{m!}{x_{1}!\cdots x_{p-1}!\big(m-\bm{1}^{\top}\bm{x}\big)!}}_{=\beta(\bm{x})}\exp\Big(\underbrace{\sum_{j\in[p-1]}x_{j}\ln\frac{\theta_{j}}{1-\bm{1}^{\top}\bm{\theta}}}_{=\bm{\gamma}(\bm{x})^{\top}\bm{\eta}(\bm{\theta})}\underbrace{+m\ln\big(1-\bm{1}^{\top}\bm{\theta}\big)}_{=-A(\bm{\theta})}\Big).$
The next proposition shows a function $G(\cdot)$ that satisfies (2.4) for (4.28) under (4.29).
@@
Let $m\in\mathbb{N}$ satisfy $m\geq\max\{\sum_{j\in[p]}n_{i,j}|i\in[I]\}$. If $G$ is defined by
$G(\bm{x})$ $=K-\sum_{i\in[I]}a_{i}g_{i}(\bm{x}),$
where $g_{i}(\cdot)$ is defined with $m$ as
$g_{i}(\bm{x}):=$ $\left\{\begin{array}[]{ll}\displaystyle\frac{\prod\limits_{j\in[p]}x_{j}(x_{j}-1)\cdots(x_{j}-n_{i,j}+1)}{m(m-1)\cdots\Big(m-\sum\limits_{j\in[p]}n_{i,j}+1\Big)},&\mbox{if }\sum\limits_{j\in[p]}n_{i,j}>0,\\ 1,&\mbox{if }\sum\limits_{j\in[p]}n_{i,j}=0,\end{array}\right$.
where $\bm{x}=(x_{1},\ldots,x_{p-1})^{\top}$ and $x_{p}=m-\bm{1}^{\top}\bm{x}$, the condition (2.4) for (4.28) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (4.29) and $\Theta=\{\bm{\theta}\in\mathbb{R}_{+}^{p-1}|\bm{1}^{\top}\bm{\theta}\leq 1\}$.
See Appendix A.8 for the proof.
Similarly to the optimization over a rectangle discussed in the previous subsection, the surrogate function is given as
$u(\bm{\theta}|\bm{\theta}^{(t-1)})=-\sum\limits_{j\in[p-1]}\Big\{\mathbb{E}_{q_{\bm{x}}(\theta^{(t-1)}_{j})}[X_{j}]\ln\frac{\theta_{j}}{\theta^{(t-1)}_{j}}\Big\}$
$\qquad\qquad\qquad\quad-\Big(m-\sum\limits_{j\in[p-1]}\mathbb{E}_{q_{\bm{x}}(\theta^{(t-1)}_{j})}[X_{j}]\Big)\ln\frac{1-\bm{1}^{\top}\bm{\theta}}{1-\bm{1}^{\top}\bm{\theta}^{(t-1)}}+f(\bm{\theta}^{(t-1)}),$
and its minimization (3.6) can be reduced to a natural gradient descent (3.9) as follows.
@@
If the multinomial distribution (4.29) is employed for Algorithm 1 to solve (4.28), the update formula (3.9) results in
$\begin{split}V^{(t-1)}&\leftarrow\sum_{j\in[p-1]}\theta_{j}^{(t-1)}\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)}),\\ \theta_{j}^{(t)}&\leftarrow\theta_{j}^{(t-1)}-\frac{\theta^{(t-1)}_{j}}{m}\Big(\frac{\partial}{\partial\theta_{j}}f(\bm{\theta}^{(t-1)})-V^{(t-1)}\Big),\quad j\in[p-1].\end{split}$ (4.30)
Similarly to (4.19), the update formula (4.30) is component-wise; no line search is involved and the use of $m=\max\{\sum_{j\in[p]}n_{i,j}|i\in[I]\}$ yields the largest step size. See Appendix A.10 for the proof. Algorithm 4 shows the full description of the procedure.
Algorithm 4 EM algorithm for minimizing the polynomial objective over a simplex (4.28)
1:Let $K>\max\{F(\bm{\lambda})|\bm{1}^{\top}\bm{\lambda}=1,\bm{\lambda}\geq\bm{0}\}$ and $\bm{\theta}^{(0)}\in\{\bm{\theta}\in\mathbb{R}_{++}^{p-1}|\bm{1}^{\top}\bm{\theta}<1\}$, and set $t\leftarrow 1$.
@@
3:  Update the solution by (4.30) and $t\leftarrow t+1$.
@@
4.4. Global convergence
Global convergence of Algorithms 3 and 4 can be shown in a unified manner. Assume that the feasible set of (2.1) is represented by a system of inequalities as $\Theta=\{\bm{\theta}\in\mathbb{R}^{p}|g_{i}(\bm{\theta})\leq 0,i\in[M]\}$ and has an interior point, i.e., $\exists\bm{\theta}$ such that $g_{i}(\bm{\theta})<0,i\in[M]$:
$\begin{array}[]{|lll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&f(\bm{\theta})\\ \mbox{subject to}&g_{i}(\bm{\theta})\leq 0,\quad i\in[M],\end{array}$ (4.33)